OpenStudy (anonymous):

y = x^4 + 2e^x; find y'

OpenStudy (anonymous):

y=x^4+2e^x

OpenStudy (anonymous):

yeah so what's the deal with e what's the derivative of e and what's the derivative of e^x

OpenStudy (anonymous):

how does all that work

OpenStudy (anonymous):

actually it is \[4x^2+2e^x\]

OpenStudy (anonymous):

\[\frac{d}{dx}e^x=e^x\]

OpenStudy (anonymous):

why 4x^2 i think that is 4x^3- Satellite pls confirm

OpenStudy (anonymous):

y=4x^3+2e^x

OpenStudy (anonymous):

yeah what cornitodisc said

OpenStudy (anonymous):